a company consumes 12000 units of a raw material. the company has a production capacity of 60 units/day, the cost of each unit produced by the company is $8. the setup and tooling up cost is $96 per set up. the capacity charges are 15% of cost per uit. Determine how frequently the production runs be made


1
Expert's answer
2021-10-22T07:05:32-0400

1200/60=20

1200×8= 9600

= 9600×0.15= 1440units